In his Chemistry class, Darin found a solution of hydrochloric acid that needed to be kept between 20 and 35 C. What is the range in temperature in degrees Fahrenheit if the Celsius/Fahrenheit conversion formula is
C= 5/9 (F-32)

Answer by malakumar_kos@yahoo.com(315)


In his Chemistry class, Darin found a solution of hydrochloric acid that needed to be kept between 20 and 35 C. What is the range in temperature in degrees Fahrenheit if the Celsius/Fahrenheit conversion formula is
C= 5/9 (F-32)

Fahrenheit/celsius formula is F = 9/5.c+32
when c = 20, F = 9/5.20+32 = 36+32 = 68
when c = 35 , F = 9/5.35+32 = 63+32 = 95
The range in temprature in degrees Fahrenheit is 68F to 95F